# Invariant sets for homeomorphisms of hyperbolic 3-manifolds

### v1

@@ 3.3. Topological entropy
Consider a small compact connected set $\mathcal{C}$ in $\widetilde{M}$ of diameter $\varepsilon>0$ whose projection to $\widetilde{\Lambda_{f}}$ contains a connected set of positive diameter $\delta>0$ within a strong unstable leaf (see Lemma 3.9). Using the uniform expansion of $\widetilde{G_{t}}$ in $T^{1}\widetilde{M}$ and the fact that orbits of $\widetilde{f}$ uniformly shadow the orbits of the geodesic flow, it follows that there is some $N>0$ (depending only on $\delta$) so that $\mathcal{C}$ contains two compact connected subsets $\mathcal{C}_{1},\mathcal{C}_{2}\subset\mathcal{C}$ so that $d(f^{N}(\mathcal{C}_{1}),f^{N}(\mathcal{C}_{2}))>\epsilon$ and the projections of $f^{N}(\mathcal{C}_{1})$ and $f^{N}(\mathcal{C}_{2})$ to $\widetilde{\Lambda_{f}}$ each contain a connected set of positive diameter $\delta>0$ within a strong unstable leaf. Repeating this procedure, we get that we can find at least $2^{k}$ points in $\mathcal{C}$ whose orbits separate more than $\varepsilon$ in some iterate between 00 and $kN$. This shows that the entropy of $f$ is at least $\frac{1}{N}\log(2)>0$.
@@ 4. Proof of Theorem B
Let $T:X\to X$ be a minimal homeomorphism and $\mathcal{A}:X\to\mathrm{PSL}_{2}(\mathbb{C})$ be [^5] In [BPS] they work with real matrices, but complex matrices can be thought as inside real matrices in the double of the dimension. Similarly, the fact that we quotient by ±$\pm\mathrm{id}$ is not an issue since the norm is still well defined. a continuous function and denote by $\mathcal{A}^{(n)}(x)=\mathcal{A}(T^{n-1}x)\cdots\mathcal{A}(x)$. Then either the set of points on which $\liminf_{n}\frac{1}{n}\log\|\mathcal{A}^{(n)}(x)\|=0$ is $G_{\delta}$-dense, or, there are constants $C>0,\tau>0$ such that for every $x\in X$ and $n>0$ one has that
@@
[…]
@@
Given a homeomorphism $f:M\to M$ of a closed hyperbolic 3-manifold and $\widetilde{f}$ a good lift, we say that a point $x\in M$ has a $\lambda$-quasi-geodesic orbit if there exists $\lambda>0$ such that $\lambda^{-1}|n-m|-\lambda\leq d(\widetilde{f}^{n}(\tilde{x}),\widetilde{f}^{m}(\tilde{x}))\leq\lambda|n-m|+\lambda$ for every $n,m\in\mathbb{Z}$ and $\tilde{x}$ lift of $x$ (i.e. the map $\mathbb{Z}\to\widetilde{M}=\mathbb{H}^{3}$ given by $n\mapsto\tilde{f}^{n}(\tilde{x})$ is a quasi-isometry).
@@ 5.2. Some properties of the pseudo-Anosov regulating flow
Figure 4. A stable line through a $3$-prong.
@@
The last property can also be phrased in terms of cross ratios of cuadruples of points in the circle at infinity. The fact that the cross ratios are far from 00 and $\infty$ is a property invariant under quasi-isometry and gives an equivalent notion. We chose this definition in order to argue in $\mathbb{H}^{2}$ where many classical results can be quoted directly.
@@ 5.2.1. Contraction and expansion of lines
where $Z:\widetilde{M}/\widetilde{\mathcal{F}}\to\widetilde{M}/\widetilde{\mathcal{F}}$ is a structure map of $\widetilde{\mathcal{F}}$.
@@ 5.3.1. Construction of good neighborhoods
    $\mathcal{W}^{wu}(\widetilde{\delta})$ does not intersect $V^{r}$ and $\mathcal{W}^{ws}(\widetilde{\delta})$ does not intersect $U^{r}$.
@@ 5.3.2. Invariance of good neighborhoods
Let $Z:\widetilde{M}/\widetilde{\mathcal{F}}\to\widetilde{M}/\widetilde{\mathcal{F}}$ be the structure map of $\mathcal{F}$. Recall that given $x\in\widetilde{M}$, ~$\widetilde{\mathcal{F}}(x)$ denotes the leaf of $\widetilde{\mathcal{F}}$ through $x$, and we write $\widetilde{f}^{k}>Z$ to denote that ~$\widetilde{\mathcal{F}}(\widetilde{f}^{k}(x))$ lies above $Z(\widetilde{\mathcal{F}}(x))$ for all $x\in\widetilde{M}$.
@@ 5.3.3. Escape properties
[…]
@@
[…]
@@
[…]
@@ 5.3.4. Construction of invariant closed sets
[…]
@@ 5.3.5. Conclusion of the proof
Note that $R^{0}=D$. Recall that $\mathcal{W}^{wu}(\widetilde{\delta})$ does not intersect $V$ and is contained in $D\cup U$. Since $\mathcal{W}^{wu}(\widetilde{\delta})$ separates $\widetilde{M}$ into two components, one containing $V_{1}$ and the other $V_{2}$, we conclude that $C$ must intersect $\mathcal{W}^{cu}$. Therefore, $C$ intersects $D\setminus U$.
@@
[…]
@@ 5.4. Further properties of $T_{\gamma}$
Recall that $\mathcal{W}^{wu}(\widetilde{\delta})$ is contained in $D\cup U$ and avoids $V$ and $\mathcal{W}^{ws}(\widetilde{\delta})$ is contained in $D\cup V$ and avoids $U$. It follows that $R^{n}\cup V$ and $Q^{n}\cup U$ contain properly embedded surfaces avoiding respectively $U$ and $V$. Up to small perturbation, we can assume that these surfaces, which we denote as $A^{u}$ and $A^{s}$ respectively, intersect transversally and their intersection is then contained in $R^{n}\cap Q^{n}$. By the transversality $A^{u}\cap A^{s}$ must be a union of properly embedded curves in $R^{n}\cap Q^{n}$ which separates $A^{u}\cap V_{1}$ from $A^{u}\cap V_{2}$ we deduce that at least one connected component must intersect every leaf of $\widetilde{\mathcal{F}}$. This is the desired component that allows us to conclude the proof. ∎
@@
Start with a quasi-geodesic flow $\varphi_{t}:M\to M$ of a closed hyperbolic 3-manifold. For instance, one can take the suspension flow of a pseudo-Anosov homeomorphism of a surface of genus $\geq 2$. Fix a regular periodic orbit, in a small neighborhood it looks like the suspension of a hyperbolic fixed point in the plane. By considering a convenient time $t_{0}$ if we consider $f_{0}=\varphi_{t_{0}}$ one can take coordinates $(x,y,t)\in[-1,1]^{2}\times S^{1}$ around the periodic orbit so that the dynamics of $f_{0}$ in these coordinates is: $(x,y,t)=(2x,\frac{1}{2}y,t)$ (here, $t\in S^{1}$ where $S^{1}$ is considered as $\mathbb{R}/_{\mathbb{Z}}$ and we think of this as making an integer amount of turns). We will remove a small neighborhood of $x=y=0$ and change the dynamics so that the maximal invariant set in this neighborhood is a pseudo-circle. Doing this, one gets $T_{\gamma}$ to project to a pseudo-circle if $\gamma$ is the deck transformation associated to this periodic orbit. For this, consider $h:[-1,1]\times S^{1}\to[-1,1]\times S^{1}$ an embedding with the following properties:
  • $h(y,t)=(\frac{1}{2}y,t)$ in a neighborhood of the boundary of $[-1,1]\times S^{1}$.
@@
Such a construction can be found for instance in [CO]. Now, since $h$ is homotopic to the identity, we can choose $h_{s}:[-1,1]\times S^{1}\to[-1,1]\times S^{1}$ with $s\in[-1,1]$ a continuous family of embeddings so that $h_{s}(y,t)=(\frac{1}{2}y,t)$ in a neighborhood of the boundary of $[-1,1]\times S^{1}$ for all $s$, and so that $h_{0}=h$ and $h_{\pm 1}(y,t)=(\frac{1}{2}y,t)$ for all $(y,t)$. Finally, we can modify $f_{0}$ to $f_{1}$ which coincides with $f_{0}$ outside this neighborhood and so that in these coordinates it is $f_{1}(x,y,t)=(2x,h_{x}(y,t))$. This has the desired properties.
@@ 6.1. Uniform positive rate and invariant measures
Note that by induction, it is enough to show that there is $T>0$ so that $Q(x)-Q(y)>T$ implies that $d(x,y)>1$. Indeed, if such a $T>0$ exists and $Q(x)-Q(y)>nT$, then one can consider $z$ in the geodesic joining $x$ and $y$ so that $Q(x)-Q(z)>T$ and $Q(z)-Q(y)>(n-1)T$, and then obtain that $d(x,y)>1+d(z,y)$. Inductively one gets that $d(x,y)>n$.
The existence of $T>0$ so that $Q(x)-Q(y)>T$ implies that $d(x,y)>1$ follows from (2.1), indeed, if there are sequences $x_{n},y_{n}$ for which $Q(x_{n})-Q(y_{n})\to\infty$ but $d(x_{n},y_{n})\leq 1$ then by compactness of $M$ we can consider deck transformations $\gamma_{n}$ which, up to taking subsequences, satisfy that $\gamma_{n}x_{n}\to x_{\infty}$ and $\gamma_{n}y_{n}\to y_{\infty}$. On the other hand, since $Q(x_{n})-Q(y_{n})\to\infty$, equation (2.1) implies that $Q(\gamma_{n}x_{n})-Q(\gamma_{n}y_{n})\to\infty$ contradicting the continuity of $Q$. This completes the proof. ∎
@@ 6.2. Homological rotation and fibers
We can consider $[S]\in H_{2}(M,\mathbb{Z})$ an integer homology class dual to $c$, by this, we mean that if $\alpha$ is a closed $1$-form with $[\alpha]=c$ then, it integrates positive on every closed curve which has positive intersection number with some representative of $S$.
@@
Recall that, given $x\in M$, we denoted $\eta_{x}^{n}$ the arc from $x$ to $f^{n}(x)$ produced by the homotopy from $\mathrm{id}$ to $f$ (cf. § 2.2). The fact that for every measure we have that $\ell_{c}(\mu)>0$ implies that there exists $n>0$ so that the signed intersection number between $\eta_{x}^{n}$ (with fixed endpoints) and $S$ is $\geq 2$ by an argument as in Proposition 2.1. If $\tilde{x}$ is a lift of $x$, we denote by $\tilde{\eta}_{x}^{n}$ to the lift of $\eta_{x}^{n}$ starting at $\tilde{x}$ (and ending at $\widetilde{f}^{n}(\tilde{x})$).
@@ 7. Non hyperbolic manifolds and some problems
Let $\phi_{t}:M\to M$ be a minimal flow on a 3-manifold with fundamental group of exponential growth and which is tangent to a two dimensional foliation. Is it true that in the universal cover $\widetilde{M}$ one has that for every $x\in\widetilde{M}$ one has that $\lim_{t}\frac{1}{t}d_{\widetilde{M}}(\widetilde{\phi_{t}}(x),x)=0$?

### v2

@@ 3.3. Topological entropy
Consider a small compact connected set $\mathcal{C}$ in $\widetilde{M}$ of diameter $\varepsilon>0$ whose projection to $\widetilde{\Lambda_{f}}$ contains a connected set of positive diameter $\delta>0$ within a strong unstable leaf (see Lemma 3.9). Using the uniform expansion of $\widetilde{G_{t}}$ in $T^{1}\widetilde{M}$ and the fact that orbits of $\widetilde{f}$ uniformly shadow the orbits of the geodesic flow, it follows that there is some $N>0$ (depending only on $\delta$) so that $\mathcal{C}$ contains two compact connected subsets $\mathcal{C}_{1},\mathcal{C}_{2}\subset\mathcal{C}$ so that $d(f^{N}(\mathcal{C}_{1}),f^{N}(\mathcal{C}_{2}))>\epsilon$ and the projections of $f^{N}(\mathcal{C}_{1})$ and $f^{N}(\mathcal{C}_{2})$ to $\widetilde{\Lambda_{f}}$ each contain a connected set of positive diameter $\delta>0$ within a strong unstable leaf. Repeating this procedure, we get that we can find at least $2^{k}$ points in $\mathcal{C}$ whose orbits separate more than $\varepsilon$ in some iterate between 0 and $kN$. This shows that the entropy of $f$ is at least $\frac{1}{N}\log(2)>0$.
We close this section by noticing that while we have exploited the sort of semiconjugacy given by the maps $e^{\pm}$ to produce information about the homeomorphism, there are still many questions which are unclear. Even in the flow case (see [FrLa]) one can produce examples which differ greatly from the models (for instance, one can blow up orbits to eliminate transitivity, or even create new dynamics by perturbation). For flows, in the recent [FrLa] it has been obtained that one can associate a natural ‘core’ dynamics to each flow, and that this core dynamics contains a pseudo-Anosov flow. In our case, we are far from obtaining something similar, except under the stronger assumptions of Theorem C where we manage to obtain a similar result.
@@ 4. Proof of Theorem B
Let $T:X\to X$ be a minimal homeomorphism and $\mathcal{A}:X\to\mathrm{PSL}_{2}(\mathbb{C})$ be [^6] In [BPS] they work with real matrices, but complex matrices can be thought as inside real matrices in the double of the dimension. Similarly, the fact that we quotient by $\pm\mathrm{id}$ is not an issue since the norm is still well defined. a continuous function and denote by $\mathcal{A}^{(n)}(x)=\mathcal{A}(T^{n-1}x)\cdots\mathcal{A}(x)$. Then either the set of points on which $\liminf_{n}\frac{1}{n}\log\|\mathcal{A}^{(n)}(x)\|=0$ is $G_{\delta}$-dense, or, there are constants $C>0,\tau>0$ such that for every $x\in X$ and $n>0$ one has that
@@
This allows one to define a linear cocycle as $\mathcal{A}:M\to\mathrm{PSL}_{2}(\mathbb{C})$ where $\mathcal{A}(x)=g_{\tilde{x}}$ where $\tilde{x}\in\widetilde{M}$ projects to $x$. Note that given $x\in\mathbb{H}^{3}$, there exists a norm $\|\cdot\|$ so that $d(\widetilde{f}^{n}(\tilde{x}),\tilde{x})=2\log\|\mathcal{A}^{(n)}(x)\|$ and therefore the dichotomy given in Theorem 4.1 translates directly in the dichotomy claimed in Corollary 4.2. ∎
@@
[…]
@@ 5.2. Some properties of the pseudo-Anosov regulating flow
Figure 4. A stable line through a 33-prong.
@@
The last property can also be phrased in terms of cross ratios of cuadruples of points in the circle at infinity. The fact that the cross ratios are far from 0 and $\infty$ is a property invariant under quasi-isometry and gives an equivalent notion. We chose this definition in order to argue in $\mathbb{H}^{2}$ where many classical results can be quoted directly.
@@ 5.2.1. Contraction and expansion of lines
where $Z:\widetilde{M}/\widetilde{\mathcal{F}}\to\widetilde{M}$$\widetilde{\mathcal{F}}$ is a structure map of $\widetilde{\mathcal{F}}$.
@@ 5.3.1. Construction of good neighborhoods
[…]
@@ 5.3.2. Invariance of good neighborhoods
Let $Z:\widetilde{M}/\widetilde{\mathcal{F}}\to\widetilde{M}$$\widetilde{\mathcal{F}}$ be the structure map of $\mathcal{F}$. Recall that given $x\in\widetilde{M}$, $\widetilde{\mathcal{F}}(x)$ denotes the leaf of $\widetilde{\mathcal{F}}$ through $x$, and we write $\widetilde{f}^{k}>Z$ to denote that $\widetilde{\mathcal{F}}(\widetilde{f}^{k}(x))$ lies above $Z(\widetilde{\mathcal{F}}(x))$ for all $x\in\widetilde{M}$.
@@ 5.3.3. Escape properties
Suppose $k$ is such that $\widetilde{f}^{k}\geq Z$, and $r_{k}$ is given by Proposition 5.15. Then, for $i=1,2$, for every $x\in U_{i}^{r_{k}}$ and $y\in V_{i}^{r_{k}}$, we have $d(\widetilde{f}^{nk}(x),\widetilde{\delta})\xrightarrow[n$$\infty$ and $d(\widetilde{f}^{-nk}(y),\widetilde{\delta})\xrightarrow[n$$\infty$.
@@
Suppose $k$ satisfies $\widetilde{f}^{k}\geq Z$, and $r_{k}$ is given by Proposition 5.15. Then, for every $x\in U^{r_{k}}$, if $L_{n}$ is the leaf through $\widetilde{f}^{nk}(x)$, we have $d_{L_{n}}(x_{L_{n}},\widetilde{f}^{nk}(x))\xrightarrow[n$$\infty$. Similarly, for every $y\in V^{r_{k}}$, if $L_{n}$ is the leaf through $\widetilde{f}^{-nk}(y)$, we have $d_{L_{n}}(x_{L_{n}},\widetilde{f}^{-nk}(y))\xrightarrow[n$$\infty$.
@@
Suppose $x\in U_{1}^{r_{k}}$. Lemma 5.10 ensures that $d_{L_{n}}(x_{L_{n}},U_{1}^{r_{k}+n}\cap L_{n})\xrightarrow[n$$\infty$, since $d_{\mathcal{G}^{u}}(x_{L_{n}},u_{L_{n}}^{r_{k}+n})=r_{k}+n\xrightarrow[n$$\infty$. Proposition 5.15 guarantees that $\widetilde{f}^{nk}(x)\in U_{1}^{r_{k}+n}$, thus $d_{L_{n}}(x_{L_{n}},\widetilde{f}^{nk}(x))\xrightarrow[n$$\infty$.
@@ 5.3.4. Construction of invariant closed sets
    For every $x\in U_{i}$, $y\in V_{i}$, it holds that $d(\widetilde{f}^{n}(x),\widetilde{\delta})\xrightarrow[n$$\infty$ and $d(\widetilde{f}^{-n}(y),\widetilde{\delta})\xrightarrow[n$$\infty$.
@@ 5.3.5. Conclusion of the proof
[…]
@@
From point 2 of Proposition 5.22, it follows that $\widetilde{M}\setminus D$ is contained in $U\cup\widetilde{f}(V)$. Thus, point 3 of Proposition 5.22 implies that every point $x$ outside $D$ satisfies $d(\widetilde{f}^{n}(x),g_{\gamma})\xrightarrow[n$$\infty$ or $d(\widetilde{f}^{-n}(x),g_{\gamma})\xrightarrow[n$$\infty$.
@@ 5.4. Further properties of $T_{\gamma}$
[…]
@@
Start with a quasi-geodesic flow $\varphi_{t}:M\to M$ of a closed hyperbolic 3-manifold. For instance, one can take the suspension flow of a pseudo-Anosov homeomorphism of a surface of genus $\geq 2$. Fix a regular periodic orbit, in a small neighborhood it looks like the suspension of a hyperbolic fixed point in the plane. By considering a convenient time $t_{0}$, if we consider $f_{0}=\varphi_{t_{0}}$ one can take coordinates $(x,y,t)\in[-1,1]^{2}\times S^{1}$ around the periodic orbit so that the dynamics of $f_{0}$ in these coordinates is: $(x,y,t)=(2x,\frac{1}{2}y,t)$ (here, $t\in S^{1}$ where $S^{1}$ is considered as $\mathbb{R}/_{\mathbb{Z}}$ and we think of this as making an integer amount of turns). We will remove a small neighborhood of $x=y=0$ and change the dynamics so that the maximal invariant set in this neighborhood is a pseudo-circle. Doing this, one gets $T_{\gamma}$ to project to a pseudo-circle if $\gamma$ is the deck transformation associated to this periodic orbit. For this, consider $h:[-1,1]\times S^{1}\to[-1,1]\times S^{1}$ an embedding with the following properties:
  • $h(y,t)=(\frac{1}{2}y,t)$ in a neighborhood of the boundary of $1,1]\times S^{1}$.
@@
Such a construction can be found for instance in [CO]. Now, since $h$ is homotopic to the identity, we can choose $h_{s}:[-1,1]\times S^{1}\to[-1,1]\times S^{1}$ with $s\in[-1,1]$ a continuous family of embeddings so that $h_{s}(y,t)=(\frac{1}{2}y,t)$ in a neighborhood of the boundary of $1,1]\times S^{1}$ for all $s$, and so that $h_{0}=h$ and $h_{\pm 1}(y,t)=(\frac{1}{2}y,t)$ for all $(y,t)$. Finally, we can modify $f_{0}$ to $f_{1}$ which coincides with $f_{0}$ outside this neighborhood and so that in these coordinates it is $f_{1}(x,y,t)=(2x,h_{x}(y,t))$. This has the desired properties.
@@ 6.1. Uniform positive rate and invariant measures
Note that by induction, it is enough to show that there is $T>0$ so that Q(x)Q(y)>TQ(x)-Q(y)>T implies that $d(x,y)>1$. Indeed, if such a $T>0$ exists and Q(x)Q(y)>nTQ(x)-Q(y)>nT, then one can consider $z$ in the geodesic joining $x$ and $y$ so that Q(x)Q(z)>TQ(x)-Q(z)>T and Q(z)Q(y)>(n1)TQ(z)-Q(y)>(n-1)T, and then obtain that $d(x,y)>1+d(z,y)$. Inductively one gets that $d(x,y)>n$.
The existence of $T>0$ so that Q(x)Q(y)>TQ(x)-Q(y)>T implies that $d(x,y)>1$ follows from (2.1), indeed, if there are sequences $x_{n},y_{n}$ for which $Q(x_{n})-Q(y_{n})\to\infty$ but $d(x_{n},y_{n})\leq 1$ then by compactness of $M$ we can consider deck transformations $\gamma_{n}$ which, up to taking subsequences, satisfy that $\gamma_{n}x_{n}\to x_{\infty}$ and $\gamma_{n}y_{n}\to y_{\infty}$. On the other hand, since $Q(x_{n})-Q(y_{n})\to\infty$, equation (2.1) implies that $Q(\gamma_{n}x_{n})-Q(\gamma_{n}y_{n})\to\infty$ contradicting the continuity of $Q$. This completes the proof. ∎
@@ 6.2. Homological rotation and fibers
We can consider $S]\in H_{2}(M,\mathbb{Z})$ an integer homology class dual to $c$, by this, we mean that if $\alpha$ is a closed $1$-form with $\alpha]=c$ then, it integrates positive on every closed curve which has positive intersection number with some representative of $S$.
@@
[…]
@@ 7. Non hyperbolic manifolds and some problems
Let $\phi_{t}:M\to M$ be a minimal flow on a 3-manifold with fundamental group of exponential growth and which is tangent to a two dimensional foliation. Is it true that in the universal cover $\widetilde{M}$ for every $x\in\widetilde{M}$ one has that $\lim_{t}\frac{1}{t}d_{\widetilde{M}}(\widetilde{\phi_{t}}(x),x)=0$?